OpenStudy (anonymous):

4x^2-4x+1/8x^3-1=??????

hartnn (hartnn):

hint : 8x^3 -1 = (2x)^3 - 1^3 = (2x-1) (4x^2+2x+1)

OpenStudy (anonymous):

im doig algebraic fractioning, but honetly i dont know how to do it,can you explain how i am supposed to get the answer???

hartnn (hartnn):

first, verify the question, secondly, whats the answer given ?

OpenStudy (anonymous):

simplify the following fraction by putting in lowest terms or by preforming the indicated operation. there arent answer options. its fill in the blank. but if i dont set i it up the way the computer wants it then it is wrong....

hartnn (hartnn):

i just asked to verify that its +4x or -2x in middle...

OpenStudy (anonymous):

+4x

hartnn (hartnn):

but in the question, you wrote -4x

OpenStudy (anonymous):

i didnt look at the qustion, you put a +4 so i put the same because i knew it was a 4 not a two. my bad. -4 is write.

hartnn (hartnn):

ok, so u got this ? 8x^3 -1 = (2x)^3 - 1^3 = (2x-1) (4x^2+2x+1)

OpenStudy (anonymous):

thank you.

hartnn (hartnn):

4x^2-4x+1= (2x-1)(2x-1) and then 2x-1 cancels out

OpenStudy (anonymous):

that was all i had to do.